OpenStudy (anonymous):

Solve for x. x2-4x+2 = -2x+5

Parth (parthkohli):

1) Add 2x to both sides. 2) Subtract 5 from both sides. 3) Solve the quadratic equation.

OpenStudy (anonymous):

x=-1,3

Parth (parthkohli):

@BougyMan Please don't give answers :) thanks!

Parth (parthkohli):

http://openstudy.com/code-of-conduct is the link you must refer.

OpenStudy (anonymous):

oh sorry about that :/ im new...

OpenStudy (anonymous):

\[x^2-4x+2=-2x+5\] \[x^2-4x+2+2x-5=0\] \[x^2-2x-3=0\] \[(x-3)(x+1)=0\]

OpenStudy (anonymous):

@ParthKohli use quadratic formula?

Parth (parthkohli):

Yeah, or factoring.

Parth (parthkohli):

@BougyMan It's fine :)

OpenStudy (anonymous):

-3 and 1?

OpenStudy (anonymous):

what does the pic have to do with it?

OpenStudy (anonymous):

@ParthKohli ?